# Searching for Accreting Compact Binaries in eROSITA DR1

### v1

@@ Abstract
Compact binaries with accreting white dwarfs, neutron stars, or black holes are crucial for understanding accretion physics. In this study, we identify accreting compact binary candidates in the eROSITA DR1 by combining Gaia astrometry and ZTF time-domain photometry. We select accreting compact binary candidates whose X-ray emission is likely accretion-driven, based on their X-ray-to-optical flux ratios and optical colors, and further identify systems with short-period ZTF light curves, thereby ruling out AGNs. The resulting 22 candidates, including two known compact binaries, represent promising candidates for spectroscopic follow-up to confirm their accreting nature and explore their physical properties. Our results demonstrate the effectiveness of combining X-ray-to-optical flux ratio and optical color, and time-domain photometry in selecting accreting compact binaries. The method is scalable and adaptable to future multi-wavelength sky surveys, offering a promising path toward a more complete census of compact binaries in the Galaxy.
@@ 1 Introduction
Compact binaries, pairing a compact object (neutron star, black hole, or white dwarf) with a companion star, are crucial for probing extreme astrophysical processes, including mass accretion and binary evolution (M. A. Abramowicz & P. C. Fragile, 2013; K. A. Postnov & L. R. Yungelson, 2014; D. Belloni & M. R. Schreiber, 2023). The detection of compact binary integrates multiple observational techniques including X-ray observations, which aim to detect those undergoing mass transfer, such as X-ray binaries and cataclysmic variables (CVs) (R. A. Remillard & J. E. McClintock, 2006; B. Warner, 1995; F. Verbunt et al., 1997). X-ray binaries have been the cornerstone of black hole searches, with discoveries relying on X-ray outbursts reaching luminosities of – erg s-1 (J. M. Corral-Santana et al., 2016). However, outburst-based searches lack the sensitivity to identify the numerous Galactic quiescent X-ray binaries (QXBs) with x-ray luminosities of – erg s-1 (J. P. Lasota, 2000). Novel approaches are therefore essential to uncover these faint systems and expand the census of compact binaries.
@@
Studies have employed X-ray observations to search for QXBs in specific regions (D. J. Helfand, 1980; C. O. Heinke et al., 2003). The Chandra Galactic Bulge Survey, led by R. I. Hynes et al. (2011), used optical spectroscopy to characterize X-ray sources, while the Swift Bulge Survey, led by A. Bahramian et al. (2021), utilized X-ray spectroscopy, variability, and optical/infrared counterparts. Despite their sophistication, these surveys cover only —a negligible fraction of the sky even when combined with XMM-Newton/Swift archival data. This highlights the critical need for all-sky surveys to detect faint QXBs.
@@
The all-sky X-ray survey by ROSAT had limited sensitivity and relatively poor positional accuracy (), hindering the identification of optical counterparts to Galactic sources (W. Voges et al., 1999; T. Boller et al., 2016). In contrast, the extended ROentgen Survey with an Imaging Telescope Array (eROSITA) survey is 25 times more sensitive than ROSAT, with significantly improved localization (P. Predehl et al., 2021). The first eROSITA All-Sky Survey (eRASS1), publicly released on 2024 January 31, contains 1 million X-ray point sources (A. Merloni et al., 2024). This enables the systematic search for QXBs across the entire Milky Way.
@@
Unlike X-ray outbursts, the luminosities of QXBs are comparable to CVs and accreting white dwarfs (U. Kolb, 1998; C. J. Hailey et al., 2018). The population of white dwarf systems overwhelmingly outnumbers neutron star and black hole systems, making them statistically dominant in any blind search for quiescent X-ray binaries (Q. Z. Liu et al., 2007; A. F. Pala et al., 2020). The primary focus of this work is the identification of QXBs in eROSITA data, while white dwarf binaries are expected to dominate the sample due to their higher abundance. For all detected compact binaries, the nature of the accretor can be classified through X-ray spectral analysis combined with multi-band photometric properties (D. A. Smith & V. S. Dhillon, 1998).
@@
In eROSITA DR1’s X-ray sources with reliable flux (, 0.2–10 ; 50 counts, S. Saeedi et al.2022), extragalactic sources dominate. The Galactic QXB population in this flux regime suffers significant contamination by active stars/binaries, which can be efficiently filtered using the X-ray-to-optical flux ratio () and color criteria established by A. C. Rodriguez (2024). This methodology has proven highly effective, with A. C. Rodriguez et al. (2025) identifying 127 CVs within 150 through precise eROSITA/Gaia DR3 astrometric matching, while X. Wang & J. Takata (2025) expanded the sample to 412 CV candidates using ZTF light-curve periodicity analysis.
@@ 2 Candidates Selection
Our goal is to identify reliable compact binary candidates in eRASS1. To this end, we require that sources in our sample satisfy the following essential criteria:
@@
These conditions ensure the selection of candidate characteristic of compact binaries while minimizing contamination from other x-ray source populations.
@@ 2.1 Counterparts
To prioritize the accuracy of counterpart identification, we adopt the following criterion:
An eRASS1 source must have exactly one optical source within a 10′′ radius.
The 10′′ matching radius is chosen based on eROSITA’s positional accuracy and statistical matching requirements. The average positional uncertainty of eRASS1 sources is typically less than 10′′, and even smaller at high confidence levels (A. Merloni et al., 2024). This radius effectively balances the need to minimize false matches while ensuring the inclusion of true optical counterparts.
The depth of an optical survey determines its ability to detect faint sources, which is critical for identifying the true optical counterparts of X-ray sources. To match the sensitivity of eROSITA, the optical catalog must have a detection limit sufficient to resolve these counterparts. Assuming an X-ray sensitivity of erg s-1cm-2 and typical QXB luminosities of –erg s-1, we estimate that eROSITA can detect such sources out to 3 kpc. At this distance, a companion star with a luminosity of 0.1–1 would have an apparent magnitude of 19–20. Consequently, we adopt Gaia DR3 as the optical catalog for counterpart identification.
@@ 2.2 Distinguishing Candidates from Active Stars
Magnetically active stars constitute a dominant class of Galactic X-ray sources (E. D. Feigelson et al., 2002; M. A. Agüeros et al., 2009; V. M. Kaspi & A. M. Beloborodov, 2017). To distinguish X-ray emission due to stellar activity from that produced by accretion, we use Gaia photometry in conjunction with the X-ray-to-optical flux ratio. For the 127,083 sources with Gaia counterparts, we retrieve , , and from Gaia DR3 (phot_g_mean_flux, phot_bp_mean_flux and phot_rp_mean_flux), along with X-ray fluxes in the 0.2–2.3 keV band from eROSITA DR1 (ML_FLUX_1). These values are plotted in the vs. BP–RP diagram shown in Figure 2.
To identify accretion-powered systems, we adopt the empirical boundary proposed by A. C. Rodriguez (2024): . Here, the optical flux is calculated as , and the X-ray flux is used directly from (ML_FLUX_1) without absorption correction. Applying this criterion yields 57,940 sources with flux ratios above the activity threshold.
Given the known discrepancies between Gaia DR3 astrometric solutions and dynamical constraints for some low mass x-ray binaries (LMXBs) (J. Wu et al., 2016; P. Gandhi et al., 2019; P. Charles et al., 2019), we do not impose cuts on astrometric parameters. As a result, the sample includes a number of extragalactic X-ray sources such as AGNs and quasars. However, these contaminants are effectively removed in the subsequent step based on short-period optical variability.
2.3 ZTF Light Curve
Photometric light curves are a powerful diagnostic for detecting periodic variability in quiescent X-ray binaries (QXBs). In such systems, tidal distortion of the companion can induce ellipsoidal modulation, while other features, such as eclipses, reflection effects, or stochastic accretion variability, may complicate or obscure the ellipsoidal modulation. Despite these complexities, time-domain analysis remains essential for identifying periodic behavior indicative of compact binaries.
@@ 2.3 ZTF Light Curve
We employ the Lomb-Scargle algorithm to search for periodic signals (N. R. Lomb, 1976; J. D. Scargle, 1981), folding the light curves with the obtained periods . The period search window is set to 0.1–4 days, with frequency boundaries defined as minimum_frequency = 0.25 and maximum_frequency = 10. For each candidate, we fold the light curves using the period of maximum power and visually inspect both the power spectrum and phase-folded light curve. This yields 324 candidates exhibiting significant periodic modulation.
@@
3 Result
We identify 22 compact binary candidates in eROSITA DR1 through a three-step selection process: (1) cross-matching to a unique Gaia optical counterpart within a 10′′ radius; (2) requiring X-ray-to-optical flux ratios exceeding the empirical threshold for stellar activity; and (3) detecting short-period variability in ZTF time-domain photometry. Among these candidates, two are previously known systems: J041920072545 (hereafter J0419) and J102347003840 (hereafter J1023). J0419 was confirmed via dynamical analysis using data from the Large Sky Area Multi-Object Fiber Spectroscopic Telescope (LAMOST) as a binary comprising an extremely low-mass pre-white dwarf and a compact object with a mass of (X.-Q. Cui et al., 2012; Z.-X. Zhang et al., 2022). J1023 exhibits a bistable accretion behavior, switching between a radio pulsar state and a LMXB phase depending on the mass transfer rate (A. M. Archibald et al., 2009). A summary of all 22 candidates is provided in Table1.
@@ 3.1 Archival Spectroscopic Survey Data Matching
We cross-match our candidates with the LAMOST and Sloan Digital Sky Survey (SDSS, R. Ahumada et al.2020). LAMOST provides time-domain spectroscopy of J0419, and reveals double-peaked emission lines indicative of an accretion disk in spectra. Additionally, LAMOST conducted low-resolution spectroscopic observations of J063213253622 (hereafter J0632), J085909053654 (hereafter J0859), and J094558292252 (hereafter J0945), which are classified as CVs in the LAMOST Low Resolution Survey. The SDSS database includes spectroscopic observations of J1023, J0859, and J0945. The spectrum of J1023 exhibits double-peaked emission lines from an accretion disk, while J0859 and J0945 show emission line features characteristic of CVs. No spectroscopic data from either LAMOST or SDSS are available for the remaining 17 candidates.
@@ 3.2 SIMBAD Classifications of Candidates
We searched for SIMBAD entries corresponding to our 22 candidates by querying their optical coordinates. SIMBAD has adopted classifications for 15 of 22 candidates, including the well-studied compact binaries J0419 and J1023. Among them, 12 classifications are based on sources from Gaia EDR3, as illustrated in Figure 3. According to Gaia EDR3, J0632, J081210040351, J0859, and J0945 are classified as CVs; J051723115359 and J052316252737 as RR Lyrae stars; and J064513195511, J073223072292, J073644131707, J080038275245, J093757171012 (hereafter J0937), and J094146060933 as eclipsing binaries. Additionally, ZTF classified J0419 and J152611102512 as BY Draconis variables. The remaining 7 candidates currently have no classification or identification recorded in SIMBAD.
@@
In the absence of sufficient photon counts for full spectral fitting of all candidates, we use fluxes from the eROSITA DR1 catalog to construct hardness ratios. The catalog provides fluxes in four energy sub-bands, with ML_FLUX_P1, ML_FLUX_P2, ML_FLUX_P3, and ML_FLUX_P4 respectively representing the 0.2 to 0.5 keV, 0.5 to 1 keV, 1 to 2 keV, and 2 to 5 keV ranges. In subsequent calculations, we denote these fluxes as P1, P2, P3, and P4. We calculate a hardness ratio (HR) for each source to characterize its spectral properties. The hardness ratio is defined as:
We then examine the distribution of HR as a function of orbital period derived from ZTF light curves (see Figure 4). Cataclysmic variables (CVs) are expected to occupy the lower-left region of this diagram, characterized by short orbital periods (8 hours) and relatively soft X-ray spectra. In contrast, neutron star or black hole binaries generally exhibit harder X-ray emission. The distribution of our 22 candidates is broadly consistent with this trend, with the lower envelope showing a positive correlation between HR and orbital period. This suggests that the combination of X-ray hardness and orbital period may provide a useful empirical tool for distinguishing CVs from more compact accretors such as neutron stars and black holes.

### v2

# Searching for Accreting Compact Object Binaries in SRG/eROSITA eRASS1

@@ Abstract
Compact object binaries with accreting white dwarfs, neutron stars, or black holes are crucial for understanding accretion physics. In this study, we identify accreting compact object binary candidates in the SRG/eROSITA eRASS1 by combining their X-ray fluxes with Gaia photometry and ZTF time-domain observations. Candidates are selected based on their location in the ”X-ray Main Sequence”, a diagram incorporating their X-ray-to-optical flux ratios and optical colors, which suggest accretion-driven X-ray emission. We identify 22 candidates in eRASS1 catalog using a three-step selection process: (1) cross-matching to a unique Gaia optical counterpart within a 10′′ radius; (2) requiring X-ray-to-optical flux ratios exceeding the ”X-ray Main Sequence”; and (3) detecting short-period variability in ZTF time-domain photometry. The resulting 22 candidates, including two previously confirmed compact object binaries, represent promising candidates for spectroscopic follow-up to confirm their accreting nature. Our results demonstrate the effectiveness of combining X-ray-to-optical flux ratios and optical colors jointly with time-domain photometry to uncover accreting compact object binaries. The approach is scalable and adaptable to future multi-wavelength sky surveys, offering a promising path toward a more complete census of compact object binaries in the Galaxy.
@@ 1 Introduction
Compact object binaries, pairing a compact object (white dwarf, neutron star, or black hole) with a companion star, are crucial for probing extreme astrophysical processes, including mass accretion and binary evolution (M. A. Abramowicz & P. C. Fragile, 2013; K. A. Postnov & L. R. Yungelson, 2014; D. Belloni & M. R. Schreiber, 2023). The detection of compact object binary integrates multiple observational techniques including X-ray observations, which aim to detect those undergoing mass transfer, such as X-ray binaries and cataclysmic variables (CVs) (R. A. Remillard & J. E. McClintock, 2006; B. Warner, 1995; F. Verbunt et al., 1997). X-ray binaries have been the cornerstone of black hole searches, with discoveries relying on X-ray outbursts reaching luminosities of – erg s-1 (J. M. Corral-Santana et al., 2016). However, outburst-based searches lack the sensitivity to identify the numerous Galactic quiescent X-ray binaries (QXBs) with X-ray luminosities of – erg s-1 (J. P. Lasota, 2000). Novel approaches are therefore essential to uncover these faint systems and expand the census of compact object binaries.
@@
Studies have employed X-ray observations to search for QXBs in specific regions (D. J. Helfand, 1980; C. O. Heinke et al., 2003). The Chandra Galactic Bulge Survey, led by R. I. Hynes et al. (2011), used optical spectroscopy to characterize X-ray sources, while the Swift Bulge Survey, led by A. Bahramian et al. (2021), utilized X-ray spectroscopy, variability, and optical/infrared counterparts. Moreover, the Galactic Bulge is a crowded region, complicating the possibility of optical spectroscopy to classify the nature of X-ray sources located there. Despite their sophistication, these surveys cover only —a negligible fraction of the sky even when combined with XMM-Newton/Swift archival data. This highlights the critical need for all-sky surveys to detect faint QXBs.
@@
ROSAT’s all-sky survey delivered key data for identifying and characterizing X-ray binaries, enabling precise studies of their properties and evolution. However, the survey had limited sensitivity and relatively poor positional accuracy (), hindering the identification of optical counterparts to Galactic sources (W. Voges et al., 1999; T. Boller et al., 2016). In contrast, the extended ROentgen Survey with an Imaging Telescope Array (eROSITA) survey is approximately 25 times more sensitive than ROSAT in the soft X-ray band, and also provides significantly improved localization accuracy of about (P. Predehl et al., 2021). The first eROSITA All-Sky Survey (eRASS1), publicly released on 2024 January 31, contains about one million X-ray point sources (A. Merloni et al., 2024). This enables the systematic search for QXBs across the entire western Milky Way, covered by this data release. The eastern Galactic hemisphere, owned by the Russian consortium, is not yet public.
@@
Unlike X-ray outbursts, the luminosities of QXBs are comparable to CVs and accreting white dwarfs (U. Kolb, 1998; C. J. Hailey et al., 2018). The population of white dwarf systems overwhelmingly outnumbers neutron star and black hole systems, making them statistically dominant in any blind search for quiescent X-ray binaries (Q. Z. Liu et al., 2007; A. F. Pala et al., 2020). The primary focus of this work is the identification of QXBs in eROSITA data, while white dwarf binaries are expected to dominate the sample due to their higher abundance. For all detected compact object binaries, the nature of the accretor can be classified through X-ray spectral analysis combined with multi-band photometric properties (D. A. Smith & V. S. Dhillon, 1998).
@@
In eROSITA DR1’s X-ray sources with reliable flux (, 0.2–10 ; 50 counts, S. Saeedi et al.2022), extragalactic sources dominate. The Galactic QXB population in this flux regime suffers significant contamination by active stars/binaries, which can be efficiently filtered using the X-ray-to-optical flux ratio () and color criteria established by A. C. Rodriguez (2024). This methodology has proven highly effective, with A. C. Rodriguez et al. (2025) identifying 23 CVs within 150 pc and a total of 590 candidate CVs within 1 kpc through precise eROSITA/Gaia DR3 astrometric matching. X. Wang & J. Takata (2025) identified 444 candidate CVs and detected possible orbital modulation in 56 sources through photometric analysis of ZTF/TESS light curves. These works rely on Gaia astrometry for candidate selection, which naturally favors nearby and optically bright systems. As a result, they provide limited coverage of accreting compact object binaries. Effective photometry-only selection methods for QXBs are still under development.
In this work, we present a multi-wavelength framework to identify accreting compact object binaries in eROSITA DR1 through combined Gaia DR3 photometry and ZTF short-period light-curve analysis. Section 2 details the selection methodology, Section 3 presents the results, and Section 4 summarizes and discusses our results.
@@ 2 Candidates Selection
Our goal is to identify reliable accreting compact object binary candidates in eRASS1. To this end, we require that sources in our sample satisfy the following essential criteria:
@@
These conditions ensure the selection of candidate characteristic of accreting compact object binaries while minimizing contamination from other X-ray source populations.
@@ 2.1 Counterparts
To prioritize the accuracy of counterpart identification, we adopt the following criterion: an eRASS1 source must have exactly one optical source within a 10′′ radius. The 10′′ matching radius is chosen based on eROSITA’s positional accuracy and statistical matching requirements. The average positional uncertainty of eRASS1 sources is typically less than 10′′, and even smaller at high confidence levels (A. Merloni et al., 2024). This radius effectively balances the need to minimize false matches while ensuring the inclusion of true optical counterparts.
The depth of an optical survey determines its ability to detect faint sources, which is critical for identifying the true optical counterparts of X-ray sources. To match the sensitivity of eROSITA, the optical catalog must have a detection limit sufficient to resolve these counterparts. Typical QXB luminosities range from to erg s-1. Using an X-ray sensitivity of erg s-1 cm-2 and a luminosity of erg s-1, we estimate a maximum detection distance of approximately 4 kpc. However, after correcting for interstellar absorption, the effective detection distance is expected to be smaller. At this distance, a companion star with a luminosity between 0.1 and 1 would have an apparent magnitude of approximately 19 to 20. Consequently, we adopt Gaia DR3 as the optical catalog for counterpart identification.
@@ 2.2 Distinguishing Candidates from Active Stars
Magnetically active stars constitute a dominant class of Galactic X-ray sources (E. D. Feigelson et al., 2002; M. A. Agüeros et al., 2009; V. M. Kaspi & A. M. Beloborodov, 2017). To distinguish X-ray emission due to stellar activity from that produced by accretion, we use Gaia photometry in conjunction with the X-ray-to-optical flux ratio. For the 127,083 sources with Gaia counterparts, we retrieve , and from Gaia DR3 (phot_g_mean_mag, phot_bp_mean_mag and phot_rp_mean_mag), along with X-ray fluxes in the 0.2–2.3 keV band from eROSITA DR1 (ML_FLUX_1). These values are plotted in the vs. diagram shown in Figure 2.
To identify accretion-powered systems, we adopt the empirical limit proposed by A. C. Rodriguez (2024), which is based on XMM-Newton 0.2-12 keV band data: . Here, the optical flux is calculated as , and the X-ray flux is used directly from (ML_FLUX_1) in 0.2-2.3 keV band. Assuming a power-law with an index of 2, our empirical boundary is approximately 1.8 times higher than that reported in A. C. Rodriguez (2024). Consequently, applying our selection criteria imposes a more stringent constraint on our sample. Applying this criterion yields 57,940 sources with flux ratios above the activity threshold.
Given the known discrepancies between Gaia DR3 astrometric solutions and dynamical constraints for some low mass X-ray binaries (LMXBs), we do not impose cuts on astrometric parameters. As a result, the sample includes a number of extragalactic X-ray sources such as AGNs and quasars. However, these contaminants are effectively removed in the subsequent step based on short-period optical variability.
2.3 ZTF Light Curves
Photometric light curves are a powerful diagnostic for detecting periodic variability in QXBs. In such systems, tidal distortion of the companion can induce ellipsoidal modulation, while other features, such as eclipses, reflection effects, or stochastic accretion variability, may complicate or obscure the ellipsoidal modulation. Despite these complexities, time-domain analysis remains essential for identifying periodic behavior indicative of accreting compact object binaries.
@@ 2.3 ZTF Light Curves
We employ the Lomb-Scargle algorithm to search for periodic signals (N. R. Lomb, 1976; J. D. Scargle, 1981), folding the light curves with the obtained periods . The period search window is set to 0.1–4 days, with frequency boundaries defined as minimum_frequency = 0.25 and maximum_frequency = 10. This choice is motivated by our focus on X-ray binaries, which generally have longer periods than two hours. For each candidate, we fold the light curves using the period of maximum power and visually inspect both the power spectrum and phase-folded light curve. This yields 324 candidates exhibiting significant periodic modulation.
@@
3 Results
We identify 22 accreting compact object binary candidates in eROSITA DR1 through a three-step selection process: (1) cross-matching to a unique Gaia optical counterpart within a 10′′ radius; (2) requiring X-ray-to-optical flux ratios exceeding the empirical threshold for stellar activity; and (3) detecting short-period variability in ZTF time-domain photometry. Among these candidates, two are previously known systems: J041920072545 (hereafter J0419) and J102347003840 (hereafter J1023). J0419 was confirmed via dynamical analysis using data from the Large Sky Area Multi-Object Fiber Spectroscopic Telescope (LAMOST, X.-Q. Cui et al. 2012) as a binary comprising an extremely low-mass pre-white dwarf and a compact object with a mass of (Z.-X. Zhang et al., 2022). J1023 is a transitional millisecond pulsar, exhibiting a bistable accretion behavior, switching between a radio pulsar state and a LMXB phase depending on the mass transfer rate (A. M. Archibald et al., 2009). A summary of all 22 candidates is provided in Table 1.
@@ 3.1 Archival Spectroscopic Survey Data Matching
We cross-match our candidates with the LAMOST and Sloan Digital Sky Survey (SDSS, R. Ahumada et al. 2020). LAMOST provides time-domain spectroscopy of J0419, and reveals double-peaked emission lines indicative of an accretion disk in spectra. Additionally, LAMOST conducted low-resolution spectroscopic observations of J063213253622 (hereafter J0632), J085909053654 (hereafter J0859), and J094558292252 (hereafter J0945), which are classified as CVs in the LAMOST Low Resolution Survey. The SDSS database includes spectroscopic observations of J1023, J0859, and J0945. As shown in Figure 3, the spectrum of J1023 exhibits double-peaked emission lines from an accretion disk, while J0859 and J0945 show emission line features characteristic of CVs. Figure 4 displays these five sources as four blue circles and one red circle, which generally correspond to the uppermost sources in the diagram, thereby indicating that they are located furthest from the X-ray main sequence towards the upper-left region in Figure 2. No spectroscopic data from either LAMOST or SDSS are available for the remaining 17 candidates. The catalog of candidate CVs out to 1000 pc by A. C. Rodriguez et al. (2025) includes four systems in our sample: J0419, J0632, J0859, and J0945. The remaining systems are not included.
@@ 3.2 SIMBAD Classifications of Candidates
We searched for SIMBAD entries corresponding to our 22 candidates by querying their optical coordinates. SIMBAD has adopted classifications for 15 of 22 candidates, including the well-studied accreting compact object binaries J0419 and J1023. Among them, 12 classifications are based on sources from Gaia EDR3, as illustrated in Figure 4. According to Gaia EDR3, J0632, J081210040351, J0859, and J0945 are classified as CVs; J051723115359 and J052316252737 as RR Lyrae stars; and J064513195511, J073223072292, J073644131707, J080038275245, J093757171012 (hereafter J0937), and J094146060933 as eclipsing binaries. Additionally, ZTF classified J0419 and J152611102512 as BY Draconis variables. The remaining 7 candidates currently have no classification or identification recorded in SIMBAD.
@@
In the absence of sufficient photon counts for full spectral fitting of 16 candidates, we use fluxes from the eROSITA DR1 catalog to construct hardness ratios. Source count rates are retrieved in four energy bands, with ML_RATE_P1, ML_RATE_P2, ML_RATE_P3, and ML_RATE_P4 respectively representing the 0.2 to 0.5 keV, 0.5 to 1 keV, 1 to 2 keV, and 2 to 5 keV ranges. In subsequent calculations, we denote these rates as P1, P2, P3, and P4. We calculate a hardness ratio (HR) for each source to characterize its spectral properties. The hardness ratio is defined as:
We then give the distribution of HR as a function of orbital period in Figure 5.
To compare our candidates with established CV and LMXB samples, we performed two additional cross-matching analyses. First, we cross-matched the eRASS1 with the catalog of 339 known LMXBs compiled by F. Fortin et al. (2024), and selected systems that (1) have measured orbital periods and (2) have companion stars classified as stellar spectral types (i.e., non-degenerate stars). Specifically, we excluded systems with white dwarf companions and retained only those with main-sequence-like donors. This yielded a reference set of 9 LMXBs with main-sequence companions (hereafter MS-LMXBs for convenience). Second, we cross-matched the eRASS1 with the CV catalog from R. A. Downes et al. (2001) and selected those with available orbital period measurements, resulting in 119 CVs. As shown in Figure 5, our candidates (blue circles) are plotted alongside comparison samples of MS-LMXBs (red squares) and CVs (green open circles).
Compared to CVs, which usually have orbital periods below 10 hours, LMXBs can extend to significantly longer periods. C. Motch et al. (1998) ’s analysis of X-ray sources in the ROSAT survey indicates that CVs exhibit systematically softer X-ray emission than neutron star or black hole systems. CVs are expected to occupy the lower-left region of Figure 5, characterized by short orbital periods and relatively soft X-ray emissions. In contrast, neutron star or black hole binaries generally exhibit harder X-ray emission. The distribution of our 22 candidates is broadly consistent with this trend, with the lower envelope showing a positive correlation between HR and orbital period. This suggests that the combination of X-ray hardness and orbital period may provide a useful empirical tool for distinguishing CVs from more compact accretors such as neutron stars and black holes.